# Which Optimizer Works Best for Physics-Informed Neural Networks and Kolmogorov-Arnold Networks?

### v1

@@ Abstract
Physics-Informed Neural Networks (PINNs) have revolutionized the computation of PDE solutions by integrating partial differential equations (PDEs) into the neural network’s training process as soft constraints, becoming an important component of the scientific machine learning (SciML) ecosystem. In its current implementation, PINNs are mainly optimized using first-order methods like Adam, as well as quasi-Newton methods such as BFGS and its low-memory variant, L-BFGS. However, these optimizers often struggle with highly non-linear and non-convex loss landscapes, leading to challenges such as slow convergence, local minima entrapment, and (non)degenerate saddle points. In this study, we investigate the performance of Self-Scaled Broyden (SSBroyden) methods and other advanced quasi-Newton schemes, including BFGS and L-BFGS with different line search strategies approaches. These methods dynamically rescale updates based on historical gradient information, thus enhancing training efficiency and accuracy. We systematically compare these optimizers on key challenging linear, stiff, multi-scale and non-linear PDEs benchmarks, including the Burgers, Allen-Cahn, Kuramoto-Sivashinsky, and Ginzburg-Landau equations, and extend our study to Physics-Informed Kolmogorov-Arnold Networks (PIKANs) representation. Our findings provide insights into the effectiveness of second-order optimization strategies in improving the convergence and accurate generalization of PINNs for complex PDEs by orders of magnitude compared to the state-of-the-art.
@@ 1 Introduction
This section begins by providing the necessary background and motivation for our study, highlighting the key challenges and objectives. Following this, we present an overview of the BFGS (Broyden–Fletcher–Goldfarb–Shanno) and SSBroyden (Self-Scaled Broyden) optimization methods, discussing their theoretical foundations and main characteristics. Section 2 presents a comprehensive study comparing different optimization methods with various line search and trust-region strategies. Specifically, in Subsections 2.1.1 and 2.1.2, we discuss a wide range of scenarios for the Burgers equation. We extend our study to more complex PDEs, including the Allen-Cahn equation in Section 2.3, the Kuramoto-Sivashinsky equation in Subsection 2.4, and the Ginzburg-Landau equation in Section 2.5. The conclusions of our study are summarized in Section 3. Finally, in Appendix A 4, we review the performance of optimizers for the Lorenz system. Additionally, in Appendix B 5, we analyze the performance of BFGS and SSBroyden in finding the global minimum of the multi-dimensional Rosenbrock function as a pedagogical example.
Figure 1: Loss landscape of $\mathcal{L}=\mathcal{L}_{\text{Data}}+\lambda\mathcal{L}_{\text{PDE}}$ for the viscous Burgers equation (15). In the left subfigure, the landscape is shown for $\lambda=0$. This loss landscape is smooth and exhibits well-defined convexity. However, it only fits the data (initial and boundary conditions) and does not converge to the solution within the domain. In contrast, the right subfigure illustrates the loss landscape for $\lambda=1.0$. Unlike the smooth landscape in the left subfigure, this landscape contains numerous local minima and saddle points, potentially leading to (non)-degenerate saddle points. The loss landscape is computed by using the method proposed by [1].
@@ 1.1 Background and Motivation
Physics-Informed Neural Networks (PINNs), introduced in 2017, are a groundbreaking development in scientific machine learning (SciML) [2, 3]. By seamlessly integrating the fundamental physical principles of a system with neural networks, PINNs offer a versatile and mesh-free framework for solving nonlinear partial differential equations (PDEs). Unlike traditional numerical methods, PINNs directly incorporate initial and boundary conditions as well as PDE residuals into their loss functions, enabling them to address both forward problems (predicting solutions) and inverse problems (e.g.,estimating unknown parameters or unknown functions). Their adaptability and scalability make PINNs particularly well-suited for tackling challenges in high-dimensional spaces and complex geometries that conventional methods struggle to handle. The network parameters are updated during training to minimize the loss function, resulting in a solution that meets the constraints applied in the loss function.
Since their introduction, PINNs have undergone numerous extensions and adaptations to enhance their applicability and address limitations in the original framework. These advancements include uncertainty quantification [4], domain decomposition techniques [5, 6, 7, 8], and the incorporation of alternative network architectures such as convolutional neural networks (CNNs) [9] and recurrent neural networks (RNNs) [10]. Innovative approaches like Generative Adversarial PINNs (GA-PINNs) [11], Physics-Informed Graph Convolutional Networks (GCNs) [12, 13], and Bayesian PINNs (B-PINNs) [14, 15, 16] have further broadened the scope of PINNs. Physics-Informed Extreme Learning Machines (PIELM) [17] combine the computational efficiency of Extreme Learning Machines (ELMs) with the physics-informed principles of PINNs. The hp-VPINN method [18] integrates variational principles with neural networks, utilizing high-order polynomial test spaces for improved accuracy.
@@
Currently, the standard optimization algorithms used in PINNs are Adam and quasi-Newton methods, such as BFGS or its low-memory variant, L-BFGS. BFGS and L-BFGS enhance significantly the performance of PINNs by incorporating second-order information of the trainable variable space to precondition it. Although these two optimizers are very commonly employed not only in PINNs but also in other optimization problems because of their theoretical and numerical properties, recent experience in PINNs has shown that other algorithms could outperform them for a great variety of problems [48, 49, 50]. Among all of them we can highlight the Self-Scaled Broyden algorithms [51], which update the approximation to the inverse Hessian matrix using a self-scaling technique. This advanced optimization technique enhances convergence by adaptively scaling gradient updates based on historical error information. Furthermore, [52] investigates self-scaling within quasi-Newton methods from the Broyden family, introducing innovative scaling schemes and updates. Urban et al. [50] further demonstrated that these advanced optimizers, when combined with rescaled loss functions, significantly enhance the efficiency and accuracy of PINNs, enabling smaller networks to solve problems with significantly greater accuracy.
@@
For example, consider a function class $\mathcal{H}_{n}$ and let $u^{*}$ be the solution to the underlying PDE. Let $m$ represent the number of training data points. The function $h_{m}$ is the minimizer of the loss with $m$ data points, while $\hat{h}$ is the function in $\mathcal{H}_{n}$ that minimizes the loss with infinitely many data points. The error between $h_{m}$ and $u^{*}$ represents the approximation error, while the error between $\hat{h}$ and $u^{*}$ is the estimation error. In practice, $\tilde{h}_{m}$ denotes the approximation obtained after a finite number of training iterations, such as the result of 1 million iterations of gradient-based optimization. The error between $\tilde{h}_{m}$ and $\hat{h}$ is referred to as the optimization error. While the approximation error is well understood, the optimization error remains a challenging area due to the highly nonlinear and nonconvex nature of the objective function. These challenges are exacerbated by the inclusion of PDE terms in the loss function, leading to degenerate and non-degenerate saddle points. As a result, optimization often requires ad hoc tricks and tedious parameter fine-tuning through trial and error.
To provide a detailed explanation, Figure 1 illustrates the loss landscape $\mathcal{L}=\mathcal{L}_{D}+\lambda\mathcal{L}_{PDE}$ for the viscous Burgers Equation (15), using the method proposed by [1]. The left subfigure of Figure 1 corresponds to $\lambda=0$, considering only the initial and boundary conditions. This loss landscape is smooth and exhibits well-defined convexity. However, it fits only the initial and boundary condition data and fails to converge to the solution within the domain. In contrast, the right subfigure depicts the loss landscape for $\lambda=1.0$. Unlike the smooth landscape on the left, this landscape is crinkled with numerous local minima and saddle points, including potentially degenerate and non-degenrate saddle points [55]. Consequently, the presence of multiple local minima can trap the optimizer, making it extremely challenging to minimize the loss function effectively. Motivated by this, in this work, we propose and analyze a suite of second-order optimizers and their applicability to linear, nonlinear, stiff, and chaotic PDEs whose solutions are computed by PINN and PIKAN types of approximation. Furthermore, we introduce a set of optimizers tailored to various classes of PDEs, achieving unprecedented convergence to machine precision with both single and double precision arithmetic.
@@ 1.2 Overview of BFGS and SSBroyden
In optimization, a fundamental challenge is determining the optimal direction and step size to transition from the current point $\mathbf{x}_{k}$ to an improved solution. Broadly, two primary approaches address this problem: line-search methods and trust-region methods. Both rely on a quadratic model to approximate the objective function around the current iterate. The key distinction lies in how they utilize this approximation, line-search methods determine the step size along a chosen direction, while trust-region methods restrict the step to a pre-defined neighborhood where the model is considered reliable.
@@ 1.2.1 Line-search methods
In this paper we explored two well-known methodologies that we briefly describe next.
Wolfe conditions:
@@
$f\left(\textbf{x}_{k}+\alpha_{k}\textbf{p}_{k}\right)$ $\leq f\left(\textbf{x}_{k}\right)+c_{1}\alpha_{k}\textbf{p}_{k}^{\top}\nabla f_{k},$ (3)
$\nabla f\left(\textbf{x}_{k}+\alpha_{k}\textbf{p}_{k}\right)^{\top}\textbf{p}_{k}$ $\geq c_{2}\nabla f\left(\textbf{x}_{k}\right)^{\top}\textbf{p}_{k},$ (4)
@@
$\left|\nabla f\left(\textbf{x}_{k}+\alpha_{k}\textbf{p}_{k}\right)^{\top}\textbf{p}_{k}\right|\leq c_{2}\left|\nabla f\left(\textbf{x}_{k}\right)^{\top}\textbf{p}_{k}\right|.$ (5)
@@ 1.2.3 Quasi-Newton methods
where $\nabla^{2}f_{k}$ is the Hessian matrix evaluated at the current iterate $\mathbf{x}_{k}$. In both cases, the search direction is then used to determine an appropriate step length $\alpha$, ensuring a sufficient decrease in $f$.
Since Newton’s search direction corresponds to the minimum of the quadratic approximation around a given point, this method is highly efficient in regions close to the solution. However, its reliance on explicitly computing the inverse Hessian matrix makes it computationally expensive for high-dimensional problems. Additionally, away from the solution, the Hessian may not be positive definite, meaning the Newton direction is not always a proper descent direction. In contrast, the steepest descent method avoids computing the Hessian or its inverse, making each iteration significantly faster than Newton’s method. Moreover, it always provides a descent direction by definition. However, its main drawback arises when converging to a solution where the Hessian exhibits a wide range of eigenvalues. In such cases, some directions cause significantly larger variations in the objective function than others, leading to an ill-conditioned optimization problem.
To balance computational efficiency with convergence performance, quasi-Newton methods have emerged as an attractive alternative. These methods approximate the Hessian matrix iteratively using only first-order derivative information, avoiding the need for explicit computation of second derivatives. Quasi-Newton methods achieve superlinear convergence by iteratively refining the Hessian approximation $\mathbf{B}_{k}$ using information from the gradients. These methods rely on update formulas, which ensure that the updated approximation satisfies the secant equation:
@@
where $\mathbf{s}_{k}=\mathbf{x}_{k+1}-\mathbf{x}_{k}$ represents the change in the iterates, and $\mathbf{y}_{k}=\nabla f_{k+1}-\nabla f_{k}$ represents the corresponding change in gradients. The specific formula to update the Hessian approximation $\mathbf{B}_{k}$ defines the different types of Quasi-Newton methods, among which we can highlight the Symmetric-Rank-One (SR1) and BFGS (Broyden, Fletcher, Goldfarb, and Shanno) methods are examples of Quasi-Newton methods. Among these methods, the BFGS approach is particularly notable for its ability to maintain symmetry and guarantee positive definiteness under mild conditions, making it one of the most widely adopted Quasi-Newton techniques.
If combined with a line-search strategy, BFGS optimization works as follows. At a given iteration $k$, the search direction is defined as
@@
where $\mathbf{H}_{k}\equiv\mathbf{B}_{k}^{-1}$. Once this direction is computed, we choose the step length $\alpha_{k}$ and then we calculate the new point from the current iterate $\mathbf{x}_{k}$ by using (2). Then, we calculate the intermediate variables $\textbf{s}_{k}$ and $\textbf{y}_{k}$ defined before, and then we perform the update in the matrix $\mathbf{B}_{k}$ as follows
@@
emonstrate the BFGS algorithm with the backtracking line search and trust-region, we present the convergence trajectory for a quadratic function defined as

### v2

@@ Abstract
Physics-Informed Neural Networks (PINNs) have revolutionized the computation of PDE solutions by integrating partial differential equations (PDEs) into the neural network’s training process as soft constraints, becoming an important component of the scientific machine learning (SciML) ecosystem. More recently, physics-informed Kolmogorv-Arnold networks (PIKANs) have also shown to be effective and comparable in accuracy with PINNs. In their current implementation, both PINNs and PIKANs are mainly optimized using first-order methods like Adam, as well as quasi-Newton methods such as BFGS and its low-memory variant, L-BFGS. However, these optimizers often struggle with highly non-linear and non-convex loss landscapes, leading to challenges such as slow convergence, local minima entrapment, and (non)degenerate saddle points. In this study, we investigate the performance of Self-Scaled BFGS (SSBFGS), Self-Scaled Broyden (SSBroyden) methods and other advanced quasi-Newton schemes, including BFGS and L-BFGS with different line search strategies approaches. These methods dynamically rescale updates based on historical gradient information, thus enhancing training efficiency and accuracy. We systematically compare these optimizers -using both PINNs and PIKANs- on key challenging linear, stiff, multi-scale and non-linear PDEs, including the Burgers, Allen-Cahn, Kuramoto-Sivashinsky, and Ginzburg-Landau equations. Our findings provide state-of-the-art results with orders-of-magnitude accuracy improvements without the use of adaptive weights or any other enhancements typically employed in PINNs. More broadly, our results reveal insights into the effectiveness of second-order optimization strategies in significantly improving the convergence and accurate generalization of PINNs and PIKANs.
@@ 1 Introduction
This section begins by providing the necessary background and motivation for our study, highlighting the key challenges and objectives. Following this, we present an overview of the BFGS (Broyden–Fletcher–Goldfarb–Shanno), SSBFGS (Self-Scaled BFGS), and SSBroyden (Self-Scaled Broyden) optimization methods, discussing their theoretical foundations and main characteristics. Section 2 presents a comprehensive study comparing different optimization methods with various line search and trust-region strategies. Specifically, in Subsections 2.1.1 and 2.1.2, we discuss a wide range of scenarios for the Burgers equation. We extend our study to more complex PDEs, including the Allen-Cahn equation in Section 2.3, the Kuramoto-Sivashinsky equation in Subsection 2.4, and the Ginzburg-Landau equation in Section 2.5. The conclusions of our study are summarized in Section 3. Finally, in Appendix A A, we review the performance of optimizers for the Lorenz system. Additionally, in Appendix B LABEL:appendix:Quantitative_Comparison, we provide a quantitative analysis of BFGS and SSBroyden in minimizing the multi-dimensional Rosenbrock function, serving as a pedagogical example to illustrate the impact of dimensionality on optimization performance.
@@ 1.1 Background and Motivation
Physics-Informed Neural Networks (PINNs), introduced in 2017, are a groundbreaking development in scientific machine learning (SciML) [1, 2]. By seamlessly integrating the fundamental physical principles of a system with neural networks, PINNs offer a versatile and mesh-free framework for solving nonlinear partial differential equations (PDEs). Unlike traditional numerical methods, PINNs directly incorporate initial and boundary conditions as well as PDE residuals into their loss functions, enabling them to address both forward problems (predicting solutions) and inverse problems (e.g.,estimating unknown parameters or unknown functions). Their adaptability and scalability make PINNs particularly well-suited for tackling challenges in high-dimensional spaces and complex geometries that conventional methods struggle to handle. The network parameters are updated during training to minimize the loss function, resulting in a solution that meets the constraints applied in the loss function. More recently, physics-informed Kolmogorov-Anrold Networks (PIKANs) were introduced in [3, 4, 5] to solve PDEs, so in the present study, we aim to investigate both PINNS and PIKANs with respect to their optimization performance.
Since their introduction, PINNs have undergone numerous extensions and adaptations to enhance their applicability and address limitations in the original framework. These advancements include uncertainty quantification [6], domain decomposition techniques [7, 8, 9, 10], and the incorporation of alternative network architectures such as convolutional neural networks (CNNs) [11] and recurrent neural networks (RNNs) [12]. Innovative approaches like Generative Adversarial PINNs (GA-PINNs) [13], Physics-Informed Graph Convolutional Networks (GCNs) [14, 15], and Bayesian PINNs (B-PINNs) [16, 17, 18, 19] have further broadened the scope of PINNs. Physics-Informed Extreme Learning Machines (PIELM) [20] combine the computational efficiency of Extreme Learning Machines (ELMs) with the physics-informed principles of PINNs. The hp-VPINN method [21] integrates variational principles with neural networks, utilizing high-order polynomial test spaces for improved accuracy.
@@
Currently, the standard optimization algorithms used in PINNs are Adam and quasi-Newton methods, such as BFGS or its low-memory variant, L-BFGS. BFGS and L-BFGS enhance significantly the performance of PINNs by incorporating second-order information of the trainable variable space to precondition it. Although these two optimizers are very commonly employed not only in PINNs but also in other optimization problems because of their theoretical and numerical properties, recent experience in PINNs has shown that other algorithms could outperform them for a great variety of problems [51, 52, 53]. Among all of them we can highlight the Self-Scaled Broyden algorithms [54], which update the approximation to the inverse Hessian matrix using a self-scaling technique. This advanced optimization technique enhances convergence by adaptively scaling gradient updates based on historical error information. Furthermore, [55] investigates self-scaling within quasi-Newton methods from the Broyden family, introducing innovative scaling schemes and updates. Urban et al. [53] further demonstrated that these advanced optimizers, when combined with rescaled loss functions, significantly enhance the efficiency and accuracy of PINNs, enabling smaller networks to solve problems with significantly greater accuracy. As Physics-Informed Neural Networks (PINNs) often suffer from conflicting gradients between multiple loss components (e.g., data and physics losses), resolving these conflicts is essential for stable and accurate training. Recent work by Wang et al. [56] rigorously proves that second-order optimizers, particularly Newton-type methods like SOAP, exhibit positive gradient alignment and can effectively mitigate such conflicts, leading to improved convergence in multi-task optimization settings like PINNs.
@@
For example, consider a function class $\mathcal{H}_{n}$ and let $u^{*}$ be the solution to the underlying PDE. Let $m$ represent the number of training data points. The function $h_{m}$ is the minimizer of the loss with $m$ data points, while $\hat{h}$ is the function in $\mathcal{H}_{n}$ that minimizes the loss with infinitely many data points. The error between $h_{m}$ and $u^{*}$ represents the approximation error, while the error between $\hat{h}$ and $u^{*}$ is the estimation error. In practice, $\tilde{h}_{m}$ denotes the approximation obtained after a finite number of training iterations, such as the result of one million iterations of gradient-based optimization. The error between $\tilde{h}_{m}$ and $\hat{h}$ is referred to as the optimization error. While the approximation error is well understood, the optimization error remains a challenging area due to the highly nonlinear and nonconvex nature of the objective function. These challenges are exacerbated by the inclusion of PDE terms in the loss function, leading to degenerate and non-degenerate saddle points. As a result, optimization often requires ad hoc tricks and tedious parameter fine-tuning through trial and error.
Figure 1: Loss landscape of $\mathcal{L}=\mathcal{L}_{\text{Data}}+\lambda\mathcal{L}_{\text{PDE}}$ for the viscous Burgers equation (16). In the left subfigure, the landscape is shown for $\lambda=0$. This loss landscape is relatively smooth and exhibits well-defined convexity. However, it only fits the data (initial and boundary conditions) and does not converge to the solution within the domain. In contrast, the right subfigure illustrates the loss landscape for $\lambda=1.0$. Unlike the smooth landscape in the left subfigure, this landscape contains numerous local minima and saddle points, potentially leading to (non)-degenerate saddle points. The loss landscape is computed by using the method proposed by [59].
To provide an intuitive explanation, Figure 1 illustrates the loss landscape $\mathcal{L}=\mathcal{L}_{D}+\lambda\mathcal{L}_{PDE}$ for the viscous Burgers Equation (16), using the method proposed by [59]. The left subfigure of Figure 1 corresponds to $\lambda=0$, considering only the initial and boundary conditions. This loss landscape is smooth and exhibits well-defined convexity. However, it fits only the initial and boundary condition data and fails to converge to the solution within the domain. In contrast, the right subfigure depicts the loss landscape for $\lambda=1.0$. Unlike the smooth landscape on the left, this landscape is crinkled with numerous local minima and saddle points, including potentially degenerate and non-degenerate saddle points [60]. Consequently, the presence of multiple local minima can trap the optimizer, making it extremely challenging to minimize the loss function effectively. Motivated by this, in this work, we propose and analyze a suite of second-order optimizers and their applicability to linear, nonlinear, stiff, and chaotic PDEs whose solutions are computed by PINN and PIKAN types of approximation. Furthermore, we introduce a set of optimizers tailored to various classes of PDEs, achieving unprecedented convergence to machine precision with both single and double precision arithmetic. To the best of our knowledge, the current study is the most systematic and comprehensive investigation of how optimzers affect the peformance of both PINNs and PIKAns.
@@ 1.2 Overview of BFGS and SSBroyden
In optimization, a fundamental challenge is determining the optimal direction and step size to transition from the current point $\mathbf{x}_{k}$ to an improved solution. In general, two primary approaches address this problem: line search methods and trust region methods. Both rely on a quadratic model to approximate the objective function around the current iterate. The key distinction lies in how they utilize this approximation; line-search methods determine the step size along a chosen direction, while trust-region methods restrict the step to a pre-defined neighborhood where the model is considered reliable.
@@ 1.2.1 Line-search methods
In this paper, we explored two well-known methodologies that we briefly describe next.
Wolfe conditions
@@
$f\left(\textbf{x}_{k}+\alpha_{k}\textbf{p}_{k}\right)$ $\leq f\left(\textbf{x}_{k}\right)+c_{1}\alpha_{k}\textbf{p}_{k}^{T}\nabla f_{k},$ (3)
$\nabla f\left(\textbf{x}_{k}+\alpha_{k}\textbf{p}_{k}\right)^{T}\textbf{p}_{k}$ $\geq c_{2}\nabla f\left(\textbf{x}_{k}\right)^{T}\textbf{p}_{k},$ (4)
@@
$\left|\nabla f\left(\textbf{x}_{k}+\alpha_{k}\textbf{p}_{k}\right)^{T}\textbf{p}_{k}\right|\leq c_{2}\left|\nabla f\left(\textbf{x}_{k}\right)^{T}\textbf{p}_{k}\right|.$ (5)
@@ 1.2.3 Quasi-Newton methods
where $\mathbf{H}_{k}\approx\left(\nabla^{2}f_{k}\right)^{-1}$ is the Hessian matrix evaluated at the current iterate $\mathbf{x}_{k}$. In both cases, the search direction is then used to determine an appropriate step length $\alpha$, ensuring a sufficient decrease in $f$.
Since Newton’s search direction is derived from minimizing a local quadratic approximation of the objective function, it is highly efficient near the solution where the loss landscape is approximately convex and well-behaved. This quadratic model enables rapid convergence, often in just a few iterations, especially when the Hessian accurately captures local curvature. However, Newton’s method requires computing and inverting the Hessian matrix, which becomes computationally prohibitive in high-dimensional settings, such as those encountered in deep learning or PINNs. Moreover, when the current iterate is far from the solution, the Hessian may be indefinite or poorly conditioned, leading to search directions that are not guaranteed to be descent directions and may even increase the loss. In contrast, the steepest descent (or gradient descent) method avoids the need for second-order information by using the negative gradient as the update direction. This makes each iteration relatively inexpensive and guarantees a descent direction at every step. However, its efficiency drastically deteriorates in the presence of ill-conditioned, which arises when the Hessian has eigenvalues that differ by several orders of magnitude. In such cases, the optimization trajectory may exhibit zigzagging behavior along the narrow valleys of the loss surface, causing slow convergence. This is particularly problematic in multi-objective or physics-informed settings, where different loss terms may induce conflicting gradients and anisotropic curvature. As a result, the steepest descent method often struggles in scenarios where curvature-aware updates, like those in quasi-Newton or second-order methods are essential for efficient and stable convergence [56].
To balance computational efficiency with convergence performance, quasi-Newton methods have emerged as an attractive alternative to full Newton’s method. These approaches iteratively approximate the Hessian matrix using only first-order derivative information, thereby avoiding the explicit computation of second derivatives. Specifically, quasi-Newton methods achieve superlinear convergence by updating an approximation $\mathbf{B}_{k}$ of the Hessian matrix, where $\mathbf{H}_{k}\equiv\mathbf{B}_{k}^{-1}$, using gradient information from previous iterations. These methods rely on update formulas, which ensure that the updated approximation satisfies the secant equation:
@@
where $\mathbf{s}_{k}=\mathbf{x}_{k+1}-\mathbf{x}_{k}$ represents the change in the iterates, and $\mathbf{y}_{k}=\nabla f_{k+1}-\nabla f_{k}$ represents the corresponding change in gradients. The specific form of the update rule defines the type of quasi-Newton method employed. Notable examples include the Symmetric Rank-One (SR1) and the widely used BFGS (Broyden–Fletcher–Goldfarb–Shanno) methods. Among these, the BFGS method stands out due to its ability to maintain symmetry and ensure positive definiteness of the Hessian approximation under mild conditions. This makes BFGS one of the most robust and commonly used quasi-Newton techniques. Once the search direction is determined—typically defined as
@@
a suitable step size $\alpha_{k}$ is selected (often via a line search strategy). The next iterate is then updated as $\mathbf{x}_{k+1}=\mathbf{x}_{k}+\alpha_{k}\mathbf{p}_{k}$ followed by computing the $\textbf{s}_{k}$ and $\textbf{y}_{k}$, and finally updating the Hessian approximation using the the matrix $\mathbf{B}_{k}$ as follows
@@
To demonstrate the BFGS algorithm with the backtracking line search and trust-region, we present the convergence trajectory for a quadratic function defined as